ISS Daily Summary Report – 05/06/16

Rodent Research-3 (RR-3) Bone Densitometer Scans: Following the completion of ten bone densitometer scans earlier this week, the crew transferred the rodents from the habitats, configured Ultrasound 2, and have conducted the remaining ten bone densitometer scans with support from ground teams. This is the final of three planned sessions this week to study the molecular and physical changes to the musculoskeletal system that happen in space. Results will expand scientists’ understanding of muscle atrophy and bone loss in space while testing an antibody that has been known to prevent muscle wasting in mice on Earth.


Microbial Observatory-1 Operations: Microbial Observatory -1 sample collections scheduled this week were completed when the crew took surface and air samples using various devices in multiple locations and photographing each location prior to sample collections. The investigation monitors the types of microbes present on ISS over a one-year period. Samples are returned to Earth for further study, enabling scientists to understand the diversity of the microbial flora on the ISS and how it changes over time.


Fluid Shifts Closeout Operations in the Service Module: Following yesterday’s successful completion of Distortion Product Otoacoustic Emission (DPOAE) and Cerebral Cochlear Fluid Pressure (CCFP) tests, today crewmembers completed the Fluid Shifts experiment run by de-configuring the CCFP analyzer and the DPOAE hardware after performing a data transfer in the Service Module (SM). The Fluids Shift investigation is divided into three one-week segments: Dilution Measures, Baseline Imaging, and Baseline Imaging using the Russian Chibis Lower Body Negative Pressure (LBNP) device. The experiment measures how much fluid shifts from the lower body to the upper body, in or out of cells and blood vessels, and determines the impact these shifts have on fluid pressure in the head, changes in vision and eye structures.


NanoRack Module 9: The fourth NanoRack Module 9 experiment session was completed earlier this morning. The crew activated, deactivated, and shook the mixture tubes to facilitate the experiment. Module-9 is a collection of student research projects utilizing the NanoRacks Mixsticks. Student teams from across the United States design their own experiments using flight approved fluids and materials. The investigation consists of several science experiments flown in a NanoRacks Module on board the ISS.


Pressurized Mating Adapter (PMA1)/Node 1 Intermodule Ventilation (IMV) Reconfiguration: Following PMA1/Node 1 IMV reconfiguration on Monday, data indicated we were still getting air exchange with the Russian Segment (RS).  Yesterday the crew covered the IMV diffuser at Node 1 Overhead Starboard 4 and constructed a cone around the diffuser section of PMA1 ducting resulting in air being redirected back to the RS to limit the air exchange through the PMA.  Overnight, significant separation was observed with an increase in RS ppCO2 (peak ~3.7 mmHg) and a decline of the USOS 24-hour average. Teams will continue monitoring CO2 at multiple locations throughout ISS.


Urine Processing Assembly (UPA) Conductivity: Ground teams are monitoring recent erratic conductivity data from the UPA distillate that triggered the system to reprocess during yesterday’s run, affecting the total distillate production.  Today’s two process cycles appeared to be nominal. Teams will continue monitoring UPA performance over the weekend and evaluate potential cause for the high conductivity and its erratic signature.


Today’s Planned Activities
All activities were completed unless otherwise noted.

  • P/TV Configure additional camcorder in LAB for HD downlink
  • Food Frequency Questionnaire
  • Rodent Research (RR) Animal transfer to MSG and back
  • Rodent Research (RR)  Retrieval from MELFI
  • FLUID SHIFTS. Comm configuration for the experiment / r/g 9995
  • FLUID SHIFTS HRF Rack 1 Ultrasound 2 activation and checkout
  • Rodent Research (RR) Experiment Ops
  • Rodent Research (RR) Bone Densitometer Scan
  • FLUID SHIFTS Ultrasound 2 Data Export
  • FLUID SHIFTS. Gathering and Connecting Equipment for TV coverage
  • NANO Viewing OBT materials for MODULE-9
  • NANO MOD9 Ops, Session 4
  • Rodent Research (RR) Bone Densitometer Scan
  • Rodent Research (RR) Bone Densitometer Scan
  • FLUID SHIFTS. Connecting OCT Laptop to RS BRI and HRF PC to ППС-26 (Instead of RSE-Med) r/g 2169
  • FLUID SHIFTS. OCT Power up in SM
  • Rodent Research (RR) Camcorder Setup
  • Counter Measure System (CMS) Harmful Contaminant Measurements in SM / r/g 2174
  • Audit of Docking and Internal Transfer System (ССВП) Accessories, Kits / r/g 2166
  • FLUID SHIFTS. CCFP Activation and Settings Adjustments in RS
  • FLUID SHIFTS. TONO Hardware setup in SM
  • FLUID SHIFTS HRF 1 Ultrasound 2 Rack Deactivation
  • FLUID SHIFTS. Chibis Setup / r/g 2168
  • FLUID SHIFTS. Operator Assistance with Chibis and Gamma-1 r/g 2168
  • Rodent Research (RR) Habitat Transfer
  • FLUID SHIFTS. Gathering Data in SM, Test Subject
  • FLUID SHIFTS. Gathering Data in SM, Operator
  • T2 Treadmill, Software Update Big Picture Words Review
  • СОЖ Maintenance
  • FLUID SHIFTS. Chibis Closeout Ops / r/g 2168
  • FLUID SHIFTS. Crew Onboard Support System (КСПЭ) Hardware Deactivation and Closing Applications
  • FLUID SHIFTS. Copying Data from CCFP/DPOAE devices and their deactivation in RS
  • FLUID SHIFTS. TONO Stowage in SM
  • Preparation for SM Window Inspection and Photography r/g 2167
  • FLUID SHIFTS. OCT Power off in SM
  • SPHEROIDS. Temperature Check
  • MICROBS1 Sample Collection Part 2
  • BAR. Experiment Ops / r/g 2172
  • Rodent Research (RR) Habitat Transfer
  • FLUID SHIFTS. Disconnecting of Laptop OCT, HRF and Connecting RSE-Med r/g 2170
  • Rodent Research (RR) Experiment Ops
  • Rodent Research (RR) Bone Densitometer Scan
  • FLUID SHIFTS. Disconnecting US converter from СУБА / r/g 2171
  • Rodent Research (RR) Bone Densitometer Scan
  • Rodent Research (RR) Bone Densitometer Scan
  • FLUID SHIFTS. Restore nominal comm config
  • Rodent Research (RR) Bone Densitometer Scan
  • Rodent Research (RR) Bone Densitometer Scan
  • FLUID SHIFTS. Final Gathering
  • FLUID SHIFTS. Hardware Handover to USOS
  • Rodent Research (RR) Post-Experiment Hardware Stowage
  • FLUID SHIFTS. Reconnecting HRF2 Laptop
  • Rodent Research (RR) Experiment Ops
  • FLUID SHIFTS OCT Laptop Setup
  • FLUID SHIFTS. Hardware Stowage
  • Inspection and Photography of SM Windows 06, 07 ,08, 09, 26 / r/g 2167
  • HAM radio session from Columbus
  • MICROBS1 Sample Collection, Part 3
  • Rodent Research (RR) Habitat Transfer
  • Reconfigure LAB camcorder and Dual Canon Li-Ion Battery Charger to nominal config
  • URAGAN. Observation and photography using Photo Equipment / r/g 2173
  • P/TV Camcorder Power Off and Stowage
  • MSG power off
  • SHD. Weekly Questionnaire


Completed Task List Items

  • None


Ground Activities

All activities were completed unless otherwise noted.

  • Fluid Shifts ops support
  • RR ops support
  • Nominal ground commanding


Three-Day Look Ahead:

Saturday, 05/07: Crew off duty, housekeeping

Sunday, 05/08: Crew off duty

Monday, 05/09: Dragon OBT, CUCU/CCP checkout, RR habitat restock, Space Pup, Sprint


QUICK ISS Status – Environmental Control Group:

                              Component Status
Elektron On
Vozdukh Manual
[СКВ] 1 – SM Air Conditioner System (“SKV1”) Off
[СКВ] 2 – SM Air Conditioner System (“SKV2”) On
Carbon Dioxide Removal Assembly (CDRA) Lab Standby
Carbon Dioxide Removal Assembly (CDRA) Node 3 Operate
Major Constituent Analyzer (MCA) Lab Idle
Major Constituent Analyzer (MCA) Node 3 Operate
Oxygen Generation Assembly (OGA) Process
Urine Processing Assembly (UPA) Standby
Trace Contaminant Control System (TCCS) Lab Off
Trace Contaminant Control System (TCCS) Node 3 Full Up